7a-b=14;solve for b
Answer by math problem solving(30)

Method 1:
7a - b = 14
7a - b + b = b + 14 ......... (add b to both sides)
7a = b + 14
7a - 14 = b + 14 - 14 ......... (subtract 14 from both sides)
7a - 14 = b
So
b = 7a - 14